# Model-guided Fuzzing of Distributed Systems

### v2

@@ 3.3. Notions of Coverage
A foundational formalization of a concurrent system’s possible executions is Mazurkiewicz traces (DBLP:conf/ac/Mazurkiewicz86, ). Traces partition the executions of a concurrent system into a set of equivalence classes (traces) based on the orderings of the concurrent events in an execution. Traces are defined on a set of events and an independence relation, where the independent events are commutative; hence, their order in execution does not affect the program behavior. The executions belonging to the same trace order the dependent events in the same total order, but they can reorder the independent events. Hence, traces partition the event sequences based on their partial order on the dependent events.
@@
Our method uses the coverage of states in the system’s abstract formal model (e.g., written in TLA+ (DBLP:books/aw/Lamport2002, )) to assess the coverage of new system behaviors and guide the test executions. The states in the abstract model succinctly represent the system’s state, abstracting away uninteresting implementation-level details while keeping the essential information of relevant system states and actions. Furthermore, high coverage over the different model states requires exploring all code paths that handle the different system states. We show in the evaluation that high coverage of the system states does not degrade the structural coverage of the relevant code.
We consider the formal model of a system as a labeled transition system ${M}=\langle Q,I,A,\delta\rangle$ where $Q$ is a set of states, $I$ is the set of initial states, $A$ is the set of actions, and $\delta\subseteq Q\times A\times Q$ is a set of transitions. An action $a\in A$ is enabled at state $q\in Q$ iff $(q,a,q^{\prime})\in\delta$ for some $q^{\prime}\in Q$. A run of ${M}$ is a sequence $\rho=q_{0}\xrightarrow{a_{1}}q_{1}\ldots\xrightarrow{a_{n}}q_{n}$ where $q_{0}\in I$ and $(q_{i},a_{i},q_{i+1})\in\delta$ holds for all $i$.
Given an execution of a system $S$ with a set of events $Events$, the formal model of that system $M$, and the mapping $\phi:Events\mapsto A$, the execution $t$ covers the abstract system states $q_{0},\ldots,q_{n}$ iff the sequence of actions $\phi(e_{1}),\ldots,\phi(e_{n})$ produce the run $\rho=q_{0}\xrightarrow{\phi(e_{1})}q_{1}\ldots\xrightarrow{\phi(e_{n})}q_{n}$ of ${M}$.
@@ 4. Implementation of ModelFuzz
After each test execution, the event mapper translates the sequence of collected system events into a sequence of abstract actions in the abstract model. The mapping is closer to syntactic mapping, and it does not require an in-depth understanding of the system under test implementation. For example, event mapping for Raft matches the protocol verbs and fields in the collected messages and the TLA+ model actions (e.g., matching the AppendEntries messages with the receiver process, term, and index numbers).
@@
Our implementation (1) converts the collected messages to a standard JSON encoding and (2) maps the standardized events in JSON format to abstract actions on the model. The interface to the controlled model checker accepts a sequence of standardized events and outputs the sequence of abstract states observed in the model.
@@
Our mapper for Raft maps system events collected during the execution into three classes of actions in the TLA+ model:
@@
  • The protocol actions for processing protocol messages,
    ClientRequest, HandleRequestVoteRequest, HandleRequest- VoteResponse,
    HandleAppendEntriesRequest, HandleAppend- EntriesRequest, HandleNilAppendEntriesResponse,
    identified with protocol arguments (e.g., term and index numbers): The system events of exchanging these messages with corresponding arguments map to the actions.
Our implementation uses TLA+ (DBLP:books/aw/Lamport2002, ) models of the distributed systems for coverage feedback, and we implement the controlled model checker for the TLC explicit state model checker (DBLP:conf/charme/YuML99, ) in the TLA+ Toolbox (DBLP:journals/corr/abs-1912-10633, ).
@@ 5. Experimental Evaluation
We implement ModelFuzz to test two industrial implementations of the Raft protocol (DBLP:conf/usenix/OngaroO14, ): Etcd-raft [^3] https://github.com/etcd-io/raft and RedisRaft [^4] https://github.com/RedisLabs/redisraft along with an implementation of a parametrized version of the example system presented in Section 2 in the Coyote framework (DBLP:conf/cloud/DeligiannisGLQ21, ). Our implementation uses the TLA+ model of the Raft protocol made available by the protocol’s authors (url/raft-tla, ) and extend it [^5] https://anonymous.4open.science/r/tlc-controlled-with-benchmarks-8E36/tla-benchmarks/Raft/model/raft_enhanced.tla by modeling (i) crash and restart of the processes and (ii) snapshot operations.
We evaluate the performance of model guidance compared to pure (unguided) random testing and guided testing using structural code coverage and trace coverage information. Specifically, we evaluate the performance of ModelFuzz in terms of test coverage and bug finding ability answering the following research questions:
@@ 2. RQ2
We address RQ1 by comparing the abstract state coverage of ModelFuzz to pure random, line coverage-guided, and trace coverage-guided fuzzing strategies. We address RQ2 by evaluating the bug-finding effectiveness of different testing strategies in two measures (DBLP:conf/icse/BohmeSM22, ): the unique number of bugs found and the number of test executions to find a bug.
Statistical evaluation (DBLP:journals/stvr/ArcuriB14, ). We analyze the statistical significance of our coverage results using the Mann-Whitney U-test (Mann_1947, ). We assess ModelFuzz’s bug-finding ability compared to the other strategies using Vargha and Delaney’s $\hat{A}_{12}$ statistic (vargha2000critique, ), with $\hat{A}_{12}=0.6$ as in previous work (DBLP:conf/ccs/MengPRS23, ).
Test configuration. We run the fuzzers with an initial set of $\lvert T_{0}\rvert=20$ random test cases. For each test case that covers a new state, we create five new test cases by mutating the original test case. We multiply the number of generated test cases proportionally with the number of new states observed in the test execution. When the set $T$ of tests to explore becomes empty, we repopulate a random set of tests and repeat the cycle until the test budget is exceeded.
Test oracle. We check the correctness of test executions by checking for assertion violations, exceptions, and crashes. We also check the serializability of the operations in etcd and Redis, running Elle (DBLP:journals/pvldb/AlvaroK20, ) on the executed operation history.
@@ 5.1. Microbenchmark in Coyote
We implemented a parametrized version of the example in Section 2 in the Coyote framework (DBLP:conf/cloud/DeligiannisGLQ21, ). The implementation parametrizes the system in (i) the number of worker processes and (ii) the number of Execute task messages that need to be processed to handle a request. For (i), we generate $m$ workers that need to register to the AppMaster before AppMaster can process a client request. For (ii), we modify the processing of Execute so that the Worker divides the work into a chain of $n$ number of tasks.
The system’s possible executions involve different interleavings of the Terminate message with the chain of Execute messages sent to the Worker. We seeded a concurrency bug that occurs if Terminate is processed by the Worker just before the last Execute message while processing a client request. The bug gets harder to trigger with increasing $m$ and $n$ since it requires all $m$ workers to register to AppMaster before Request and also to deliver the chain of Execute messages except for the last one before Terminate.  [^6] Available at https://anonymous.4open.science/r/coyote-modelfuzz-5757
@@ 5.2. Etcd-raft
Etcd-raft [^7] https://github.com/etcd-io/raft powers the popular distributed key-value store etcd [^8] https://etcd.io. It is a well-tested, production-ready implementation of Raft used by companies such as Cloudflare. We instrument its source code  [^9] https://anonymous.4open.science/r/etcd-fuzzing-29D8/README.md to intercept the messages exchanged between processes and implement the fuzzer.
We tested the executions of Etcd-raft with three processes and with five client requests. We ran our tests with a crash quota of 10 and delivered a maximum of 5 messages at each step. We report the results over an average of 20 runs, each with 100k test iterations.
@@
ModelFuzz 1013
Random 599
Line 1342
Trace 940
@@
Figure LABEL:fig:etcd-cov-1 reports the test coverage of the test harnesses in the number of abstract states observed with different strategies. The results show that model-guided test generation of ModelFuzz outperforms unguided random testing (Random), coverage-guided fuzzing using line coverage (Line), and coverage-guided fuzzing using trace coverage (Trace) in the explored number of unique system states. ModelFuzz covers 1.22x more states than Random, 1.23x more states than Line, and 1.48x more states than Trace coverage approaches. Comparing model-guided fuzzing and structural guidance, we find that in both cases, the code coverage saturates at $47.9\%$. Similarly, comparing model-guided fuzzing and trace-based fuzzing, we find that in both cases, we explore 10k unique traces. To mitigate the randomness, we perform Mann-Whitney U Test and conclude that ModelFuzz’s coverage of model states is significantly higher than all other approaches, with p values of $1.53\mathrm{e}{-10}$.
@@
Etcd-raft has been the subject of many extensive testing approaches. However, we find one new bug in addition to reproducing a seeded bug. The seeded bug modifies the condition for checking if a process has a quorum of votes. Specifically, we change the valid quorum size from $n/2+1$ to $n/3+1$. The new bug we found is more subtle and leads to a process crash when accessing a missing snapshot. We reported the bug [^10] https://github.com/etcd-io/raft/issues/108 to developers.

### v3

@@ 3.3. Notions of Coverage
A foundational formalization of a concurrent system’s possible executions is Mazurkiewicz traces (Mazurkiewicz, 1986). Traces partition the executions of a concurrent system into a set of equivalence classes (traces) based on the orderings of the concurrent events in an execution. Traces are defined on a set of events and an independence relation, where the independent events are commutative; hence, their order in execution does not affect the program behavior. The executions belonging to the same trace order the dependent events in the same total order, but they can reorder the independent events. Hence, traces partition the event sequences based on their partial order on the dependent events.
@@
Our method uses the coverage of states in the system’s abstract formal model (e.g., written in TLA+ (Lamport, 2002)) to assess the coverage of new system behaviors and guide the test executions. The states in the abstract model succinctly represent the system’s state, abstracting away uninteresting implementation-level details while keeping the essential information of relevant system states and actions. Furthermore, high coverage over the different model states requires exploring all code paths that handle the different system states. We show in our evaluation that high coverage of the formal system states does not degrade the structural code coverage.
We consider the formal model of a system as a labeled transition system ${M}=\langle Q,I,A,\delta\rangle$ where $Q$ is a set of states, $I$ is the set of initial states, $A$ is the set of actions, and $\delta\subseteq Q\times A\times Q$ is a set of transitions. An action $a\in A$ is enabled at state $q\in Q$ iff $(q,a,q^{\prime})\in\delta$ for some $q^{\prime}\in Q$. A run of M{M} is a sequence $\rho=q_{0}\xrightarrow{a_{1}}q_{1}\ldots\xrightarrow{a_{n}}q_{n}$ where $q_{0}\in I$ and $(q_{i},a_{i},q_{i+1})\in\delta$ holds for all $i$.
Given an execution of a system $S$ with a set of events $Events$, the formal model of that system $M$, and the mapping $\phi:Events\mapsto A$, the execution $t$ covers the abstract system states $q_{0},\ldots,q_{n}$ iff the sequence of actions $\phi(e_{1}),\ldots,\phi(e_{n})$ produce the run $\rho=q_{0}\xrightarrow{\phi(e_{1})}q_{1}\ldots\xrightarrow{\phi(e_{n})}q_{n}$ of M{M}.
@@ 4.1. The Workflow
Controlled scheduler
The controlled scheduler controls the nondeterminism in the delivery order of messages and executes a particular schedule of events. For each event in an event schedule, $t=\left<{\mathtt{buff}_{0}:a_{0}}\right>\ldots\left<{\mathtt{buff}_{n}:a_{n}}\right$> with $a_{0},\ldots,a_{n}\in\{\mathtt{deliver},\mathtt{crash},$ $\mathtt{restart}\}$, the controlled scheduler either delivers a message to its recipient process, or crashes a process, or restarts a process. To enforce the delivery of a message at a certain time in execution, the scheduler intercepts all in-flight messages in the system and collects them in the message buffers for each process. For the process crash and restart actions, we provide scripts for starting and stopping the processes in the system under test. The controlled scheduler runs these scripts to enforce a process crash or restart.
Event mapper.
@@
The event mapper translates the sequence of concrete events executed on the system by the controlled scheduler into a sequence of actions in the abstract model. Since the implementation level events recorded by the controlled scheduler are specific to the implementation under test, they cannot directly be mapped onto the abstract actions on the formal model. ModelFuzz uses an event mapping interface that the developer implements to map the implementation-level events to the abstract model actions. The mapping of the events from the implementation to the model is mostly syntactic, and it does not require an in-depth understanding of the system under test implementation. For example, event mapping for Raft matches the protocol verbs and fields in the collected messages and the TLA+ model actions (e.g., matching the AppendEntries messages with the receiver process, term, and index numbers).
@@
Our implementation (1) converts the collected implementation-level events to a standard JSON encoding and (2) maps the standardized events in JSON format to abstract actions on the model. The second step is embedded in the controlled TLC model checker. The interface to the controlled model checker accepts a sequence of standardized events and outputs the sequence of abstract states observed in the model.
Controlled model checker
The sequence of mapped actions is run on the controlled model checker, which enforces their execution on the model. Unlike a standard model checker, which explores the whole state space of the abstract model, the controlled model checker runs a given sequence of actions and returns the visited set of model states as coverage information. The actions in the specifications used in our evaluation are deterministic, and each execution of the implementation matches a single path in the abstract model. For the models where a single action can lead to multiple model states, model coverage provides an overapproximation, e.g., tracing all paths and collecting all states that could be visited with the executed event schedule.
For the controlled model checker, we implement a simulation engine that (i) controls the next action to take at each state of the model execution and (ii) records the visited states to provide them as feedback to the fuzzer. Our implementation uses TLA+ (Lamport, 2002) models of the distributed systems for coverage feedback, and we implement the controlled model checker for the TLC explicit state model checker (Yu et al., 1999) in the TLA+ Toolbox (Kuppe et al., 2019).
4.2. Testing Distributed Systems with ModelFuzz
ModelFuzz can be used to test any distributed system that has a TLA+ model. To test an implementation of a distributed system using ModelFuzz, two key steps are required: (i) intercepting the messages of the system under test to control their delivery and enforce the delivery ordering specified by the fuzzing algorithm, and (ii) writing an event mapper to map the system’s events into the actions defined in the TLA+ model. Testing different system implementations and using different TLA+ models do not require any modifications for the controlled TLC model checker.
The first requirement for controlling delivery is a common requirement across all controlled concurrency testing algorithms. This is essential to enable the execution of specific event schedules generated by the algorithms. Additionally, ModelFuzz requires a TLA+ model of the system as well as an event mapping. TLA+ models are available for popular distributed system protocols, such as Two-Phase Commit, Paxos, and Raft. The event mapping, which maps the implementation events to the abstract TLA+ model actions, is mostly synthetic. It is tailored to the specific system being tested and is provided by the developers.
In the rest of this section, we explain our implementation for testing Raft protocol implementations, providing an overview of the protocol, its abstract model, and the event mapper.
4.2.1. Testing Implementations of the Raft Protocol
Raft protocol.
Raft (Ongaro and Ousterhout, 2014) is a leader-based consensus protocol that proceeds in a sequence of terms. Each term begins with a leader election phase. A candidate sends RequestVote messages requesting votes from other processes to be elected as leader. A candidate process transitions to the leader state if it receives a quorum of accepted votes. Upon successful election, the leader executes client requests, records the executed operations in a log, and replicates the log of operations on the other processes by sending AppendEntries messages to the follower processes. The leader also periodically sends Heartbeat messages to communicate its availability to other processes. If a process does not receive Heartbeat messages from the leader over a time period, it moves to a new term, incrementing its term number, and transitions to the candidate state.
Abstract protocol model
We use the TLA+ model of the Raft protocol made available by the protocol’s authors (Github., [n. d.]) and extend it by modeling (i) crash and restart of the processes and (ii) snapshot operations. Therefore, besides the variables for the internal states of each process (e.g., the term number, its log of requests), the extended model uses an additional variable to keep the set of active processes in a cluster and introduces crash and restart actions. The crash and restart actions are enabled for the active and crashed processes, respectively, and they update the set of active processes in the cluster. Based on empirical evidence that implementation bugs can occur in the snapshot processing logic, we also extend the model to capture snapshot operations so that the model can guide the executions toward executions where processes trigger snapshots and restore them upon recovery. While the existing TLA+ model does not capture snapshot operations, our model uses a snapshot index variable for each process to record the snapshot index of each process and introduces actions to update the processes’ snapshot indices. Our extended TLA+ model [^3] https://github.com/burcuku/tlc-controlled-with-benchmarks/blob/main/tla-benchmarks/Raft/model/raft_enhanced.tla satisfies the correctness specifications listed in the original model and, therefore, preserves correctness.
Collection of abstract states
While the abstraction provided by the TLA+ model is useful in guiding the fuzzer, we observe the need for further abstracting the set of observed states to guide the exploration to ‘interesting’ parts of the state space. Concretely, the states of the existing TLA+ model are defined by the local states (e.g., operation logs) of each process, along with the current term numbers of each process. Therefore, the model states differentiate between the system states with the same set of local process states if they are reached in different term numbers. Consider the leader election phase of an execution. Although the local states of the processes do not change, unsuccessful leader election rounds result in hitting new system states since the processes increment their term numbers. Such state information guides the fuzzer toward exploring executions with growing term numbers without covering interesting system behavior.
To guide the fuzzer with more concise state information, we abstract the term numbers in the formal model state. Specifically, we merge two consecutive states in an execution, which differ only in terms of the number of non-leaders.
Note that state abstraction does not require any modification of the formal TLA+ model or the model checker. The abstraction post-processes the collected abstract states from the model checker, and the abstracted states are communicated to the fuzzer. In general, any abstraction can be incorporated by writing an abstraction function from the output state of the model checker to the input state of ModelFuzz.
Event mapper from Raft implementations to the abstract TLA+ actions
@@ 4.2.1. Testing Implementations of the Raft Protocol
We implemented event mappers for two different implementations of the Raft protocol (in etcd and Redis) to the abstract TLA+ model. Our mapper for Raft maps system events collected during the execution into three classes of actions in the TLA+ model:
@@
  • The protocol actions for processing protocol messages, ClientRequest, HandleRequestVoteRequest, HandleRequestVoteResponse, HandleAppendEntriesRequest, HandleAppendEntriesRequest, and
    HandleNilAppendEntriesResponse, identified with protocol arguments (e.g., term and index numbers): The system events of exchanging these messages with corresponding arguments map to the actions.
@@ 5. Experimental Evaluation
We implement ModelFuzz to test two industrial implementations of the Raft protocol (Ongaro and Ousterhout, 2014): Etcd-raft [^4] https://github.com/etcd-io/raft and RedisRaft [^5] https://github.com/RedisLabs/redisraft along with an implementation of the Two-Phase Commit Protocol and an implementation of a parametrized version of the example system presented in Section 2 in the Coyote framework (Deligiannis et al., 2021). Our implementation for the Raft protocol uses the TLA+ model of the protocol made available by the protocol’s authors (Github., [n. d.]) and extends it [^6] https://github.com/burcuku/tlc-controlled-with-benchmarks/blob/main/tla-benchmarks/Raft/model/raft_enhanced.tla by modeling (i) crash and restart of the processes and (ii) snapshot operations. Our implementation for the Two-Phase Commit protocol is based on the protocol’s model available in the TLA+ documentation (Lamport, [n. d.]) and extends it [^7] https://github.com/egeberkaygulcan/2PC-Fuzzing/blob/main/tla/TPCL.tla with transaction variables and support for multiple requests.
We evaluate the performance of model guidance compared to pure (unguided) random testing, guided testing using structural code coverage and trace coverage information, and, finally, state-of-the-art reinforcement learning guided testing, BonusMaxRL (Borgarelli et al., 2024). For comparison against BonusMaxRL, we leverage the publicly available implementation [^8] https://github.com/zeu5/raft-rl-test [^9] https://github.com/zeu5/dist-rl-testing and retain the configurations detailed in the original paper.
We evaluate the performance of ModelFuzz in terms of test coverage and bug finding ability, answering the following research questions:
@@ 2. RQ2
We address RQ1 by comparing the abstract state coverage of ModelFuzz to pure random, line coverage-guided, and trace coverage-guided fuzzing strategies. We address RQ2 by evaluating the bug-finding effectiveness of different testing strategies in two measures (Böhme et al., 2022): the unique number of bugs found and the number of test executions to find a bug.
Test oracle.
We check the correctness of test executions by checking for assertion violations, exceptions, and crashes. We also check the serializability of the operations in etcd and Redis, running Elle (Alvaro and Kingsbury, 2020), a black-box isolation checker on the collected history of log operations. on the executed operation history.
Statistical evaluation (Arcuri and Briand, 2014)
. We analyze the statistical significance of the coverage results of our experiments by performing the Mann-Whitney U-test (Mann and Whitney, 1947). We assess ModelFuzz’s bug-finding ability compared to the other strategies using Vargha and Delaney’s $\hat{A}_{12}$ statistic (Vargha and Delaney, 2000), with $\hat{A}_{12}=0.6$ as in previous work (Meng et al., 2023).
Test configuration.
We run the fuzzers with an initial set of $\lvert T_{0}\rvert=20$ random test cases. For each test case that covers a new state, we create five new test cases by mutating the original test case. We multiply the number of generated test cases proportionally with the number of new states observed in the test execution. When the set $T$ of tests to explore becomes empty, we repopulate a random set of tests and repeat the cycle until the test budget is exceeded.
@@ 5.1. Microbenchmark in Coyote
We implemented a parametrized version of the example in Section 2 in the Coyote framework (Deligiannis et al., 2021). The implementation parametrizes the system in (i) the number of worker processes and (ii) the number of Execute task messages that need to be processed to handle a request. For (i), we generate $m$ workers that need to register to the AppMaster before AppMaster can process a client request. For (ii), we modify the processing of Execute so that the Worker divides the work into a chain of $n$ number of tasks.
The system’s possible executions involve different interleavings of the Terminate message with the chain of Execute messages sent to the Worker. We seeded a concurrency bug that occurs if Terminate is processed by the Worker just before the last Execute message while processing a client request. The bug gets harder to trigger with increasing $m$ and $n$ since it requires all $m$ workers to register to AppMaster before Request and also to deliver the chain of Execute messages except for the last one before Terminate.  [^10] Available at https://github.com/burcuku/coyote-concurrency-testing
@@
5.2. Two-Phase Commit
The Two-Phase Commit protocol is a popular distributed algorithm that ensures all nodes in a distributed database either commit or abort a transaction in a coordinated manner. It plays a crucial role in maintaining consistency within distributed databases and transaction management systems. We provide a thread-based implementation, where we modify the standard protocol to handle requests as sets of variables written by a transaction [^11] https://github.com/egeberkaygulcan/2PC-Fuzzing. Each node maintains a lock table, which starts in an unlocked state. When a transaction commit request is received, the node checks the lock status of the variables. If none are locked, the vote is approved; otherwise, the request is aborted.
The implementation omits details such as network communication and storage code, as the benchmark’s primary goal is to observe the protocol’s behavioral changes under different request and server interleavings. This results in a compact server implementation of 350 lines of Go code, with approximately 150 additional lines for instrumentation. The implementation of ModelFuzz, including the fuzzer loop and communication with the controlled TLC model checker, is approximately 1k lines of Go code.
We tested the Two-Phase Commit protocol using three resource managers (RM), a dedicated transaction manager, two variables (V), and five transaction commit requests (N). These parameters were selected for two key reasons: (i) smaller V values reduce the likelihood of successfully committing concurrent requests, and (ii) the number of distinct abstract model states grows exponentially with increasing N and RM. In experiments with higher N and RM values, we observed similar results, with the only difference being a larger state space requiring a greater time budget for exploration. The chosen parameter set strikes a balance between exploration difficulty and maintaining a manageable state space.
Since the protocol is not fault-tolerant, we excluded crash faults from this benchmark. Each test execution ran for 100 steps per iteration, with a maximum of five messages delivered at each step. We run all tests for one hour and report the average result of 20 test runs.
Guidance Avg. #states
ModelFuzz 17.74M
Random 15.29M
Trace 14.41M
Line 14.87M
RL 96.98K
Figure 4. (a) Test coverage of abstract states of Two-Phase Commit. (b) Average number of abstract states discovered by the strategies.
Coverage.
Figure 4(a) presents the test coverage of the test harnesses, measured by the number of abstract states observed across different strategies. The results show that ModelFuzz surpasses both random and trace coverage-guided exploration strategies. In Figure 4(b), we report the average number of distinct abstract model states discovered by each strategy, with the best result highlighted in bold. Overall, ModelFuzz observes 1.16x more states than random exploration, 1.23x more than trace coverage-guided exploration, 1.20x more than line coverage-guided exploration, and 182.96x more than the reinforcement learning-based testing BonusMaxRL.
Our analysis of the experiments reveals that BonusMaxRL’s poor performance stems from its high learning cost, which exceeds the execution cost of the program. Within the same time budget, fuzzing-based strategies complete three orders of magnitude more iterations than BonusMaxRL. As a result, BonusMaxRL is significantly less efficient at exploring the state space.
Consistent with our other benchmarks, we address  RQ1 by confirming that model guidance outperforms random exploration, reinforcement learning-guided and other coverage-guidance strategies in the coverage of explored states.
We conduct Mann-Whitney U tests to compare our benchmarks against ModelFuzz and validate the statistical significance of our results. Across 20 experiments, the tests confirm that ModelFuzz discovers significantly more states than the other testing strategies, with p-values of $6.80\mathrm{e}{-8}$ for all comparisons.
@@ 5.3. Etcd-raft
Etcd-raft [^12] https://github.com/etcd-io/raft powers the popular distributed key-value store etcd [^13] https://etcd.io. It is a well-tested, production-ready implementation of Raft used by companies such as Cloudflare. We instrument its source code  [^14] https://github.com/zeu5/raft-fuzzing to intercept the messages exchanged between processes and implement the fuzzer. The highly modular nature of the etcd implementation allows us to instrument and run all the nodes within a single OS process, making it easier and faster to start, stop, and restart all the nodes. The Etcd-raft implementation consists of approximately 7k lines of Go code. We instrument its source code and use an additional 1k lines to implement the fuzzer loop, gain control of the messages exchanged between processes, and implement necessary adapters to communicate with the model checker.
We tested the executions of Etcd-raft with three processes and with five client requests. We ran our tests with a crash quota of 10 and delivered a maximum of 5 messages at each step. For each test run, we execute the test for 12 hours, where each iteration runs for a fixed duration of 1s, and report the average results of 20 test runs.
@@
ModelFuzz 0.243
Random 0.143
Line 0.322
Trace 0.256
RL 0.234
@@
Figure LABEL:fig:etcd-cov-1 reports the test coverage of the test harnesses in the number of abstract states observed with different strategies. The results show that model-guided test generation of ModelFuzz covers similar coverage as random testing (Random), coverage-guided fuzzing using line coverage (Line), and outperforms coverage-guided fuzzing using trace coverage (Trace) in the explored number of unique system states. However, the reinforcement learning approach BonusMaxRL outperforms all other approaches.
Given the nature of our instrumentation with Etcd-raft implementation, ModelFuzz suffers from the performance overhead of communicating with the TLC model checker to measure and retrieve the state trace for the explored execution. However, other approaches are able to achieve better coverage in the absence of this overhead. Comparing model-guided fuzzing and structural guidance, we find that in both cases, the code coverage saturates at $47.9\%$. Similarly, comparing model-guided fuzzing and trace-based fuzzing, we find that in both cases, we explore 10k unique traces.
@@
Etcd-raft has been the subject of many extensive testing approaches. However, we found one new bug in addition to reproducing a seeded bug. The seeded bug modifies the condition for checking if a process has a quorum of votes. Specifically, we change the valid quorum size from $n/2+1$ to $n/3+1$. The new bug we found is more subtle and leads to a process crash when accessing a missing snapshot. We reported the bug [^15] https://github.com/etcd-io/raft/issues/108 to developers.